ON A GENERALIZATION OF SCHNHARDT’S POLYHEDRON

JRG RAMBAU

Abstract. We show that the non-convex twisted prism over an n-gon cannot be triangulated without new vertices. For this, it does not matter what the coordinates of the n-gon are as long as the top and the bottom n-gon are congruent and the twist is not too large. This generalizes Schönhardt’s polyhedron, which is the non-convex twisted prism over a triangle.